OpenStudy (anonymous):

Whitney is taking Modeling 101. On her first modeling quiz, she received a grade of 75%. On her second modeling quiz, she received an 85%. She has her upcoming final exam, and wants to know what grade she needs to receive a C (which in this class, requires a 50% average on all coursework). If quizzes all hold the same weight, and her final is 60% of her grade, what grade must she achieve on her final?

OpenStudy (lukecrayonz):

75+85/2=80 is her current grade

OpenStudy (lukecrayonz):

so umm, one sec i remember this haha

OpenStudy (lukecrayonz):

Okayyyyy I remember

OpenStudy (anonymous):

you've done this problem before?

OpenStudy (lukecrayonz):

Jesusss I am forgetting everything I know about math right now..

OpenStudy (anonymous):

NEverrrr mind. I got it. 30. [4(80) + 6x]/10 = 50 320 + 6x = 500 6x = 180 x = 30

OpenStudy (anonymous):

Thanks for trying though.

OpenStudy (lukecrayonz):

Hmm, I got like 10.8 somehow, thankfully you didn't listen to me, but thanks for showing work so I can remember everything, horrible night for math for me and its 3:11 A.M

OpenStudy (anonymous):

lol. Same here.